Find an equation of the line having the given slope and containing the given plot. m=2,(5,7)The equation of the line is y=?
Answer by stanbon(75887)

Find an equation of the line having the given slope and containing the given plot. m=2,(5,7)The equation of the line is y=?
-----------------
Form: y = mx+b
7 = 2*5 + b
b = -3
----
Equation:
y = 2x-3
=================
Cheers,
Stan H.